The Ksp of Ca(OH)2 is 6.5 x 10-6. Will a precipitate form if the pH of a 0.085 M CaCl2 solution is adjusted to 8.8?
Answer #1

Concentration of CaCl2 = 0.085 M

So, [Ca2+] = 0.085 M

Then

pH = 8.8

So, pOH = 14 - 8.8 = 5.2

[OH-] = 10-5.2 M = 6.31× 10-6 M.

Now, Q of Ca(OH)2

= [Ca2+]  × [OH-]2

= 0.085 × ( 6.31× 10-6)2

= 3.38 × 10-12 .

Now, Ksp of Ca(OH)2 = 6.5×10-6

As, Ksp > Q , no precipitate will form.